OpenStudy (anonymous):

Hi everyone! Can someone please show me steps to solve L{sintU(t-(pi/2))} ? We need to use 2nd Translation/Shifting Theorem. Thanks :o)

OpenStudy (btaylor):

What is the theorem you need to use?

OpenStudy (anonymous):

2nd Translation/Shifting Theorem

OpenStudy (btaylor):

What does that theorem say?

OpenStudy (anonymous):

one sec...my keyboard is out of batteries and I am clicking all these letters one by one...

OpenStudy (btaylor):

alright

OpenStudy (btaylor):

I looked it up and don't know it yet. Sorry.

OpenStudy (anonymous):

L{f(t-a)U(t-a)} = e^(-as)F(s)

OpenStudy (anonymous):

i know a=pi/2

OpenStudy (kainui):

I honestly don't know, but I just started watching this youtube video to see if I can figure it out https://www.youtube.com/watch?v=ns2F_WPZxvQ

OpenStudy (kainui):

Interesting at about 3 minutes in he says that if the two functions aren't shifted by the same amount you can't do it directly, but it looks like we can rewrite \(\sin(t) = \cos (t- \pi/2)\) So that looks like our first step to plug that in!

OpenStudy (anonymous):

Hi everyone! Can someone please show me steps to solve L{sintU(t-(pi/2))} ? We need to use 2nd Translation/Shifting Theorem. Thanks :o) $$\begin{align*}\mathscr{L}\left\{\sin(t)\mathscr{U}\left(t-\frac\pi2\right)\right\}&=\mathscr{L}\left\{\sin\left(\left(t-\frac\pi2\right)+\frac\pi2\right)\mathscr{U}\left(t-\frac\pi2\right)\right\}\\&=e^{-\frac\pi2s}\mathscr{L}\{\sin\left(t+\frac\pi2\right)\}\\&=e^{-\frac\pi2s}\mathscr{L}\{\cos(t)\}\\&=e^{-\frac\pi2s}\cdot\frac{s}{s^2+1}\end{align*}$$

OpenStudy (anonymous):

using the fact that \(\sin(t+\pi/2)=\cos(t)\)